# Foundation Model for Composite Materials and Microstructural Analysis

### v1

@@ Abstract
The rapid advancement of machine learning has unlocked numerous opportunities for materials science, particularly in accelerating the design and analysis of materials. However, a significant challenge lies in the scarcity and high cost of obtaining high-quality materials datasets. In other fields, such as natural language processing, foundation models pre-trained on large datasets have achieved exceptional success in transfer learning, effectively leveraging latent features to achieve high performance on tasks with limited data. Despite this progress, the concept of foundation models remains underexplored in materials science. Here, we present a foundation model specifically designed for composite materials. Our model is pre-trained on a dataset of short-fiber composites to learn robust latent features. During transfer learning, the MMAE accurately predicts homogenized stiffness, with an R2 score reaching as high as 0.959 and consistently exceeding 0.91, even when trained on limited data. These findings validate the feasibility and effectiveness of foundation models in composite materials. We anticipate extending this approach to more complex three-dimensional composite materials, polycrystalline materials, and beyond. Moreover, this framework enables high-accuracy predictions even when experimental data are scarce, paving the way for more efficient and cost-effective materials design and analysis.
keywords:
foundation models, composite materials, transfer learning, self-supervised learning, latent features
@@ 1 Introduction
The concept of foundation models has also been successfully applied to computer vision. Masked autoencoders (MAEs), introduced by He et al. [12], employ a strategy analogous to BERT by masking portions of the input image and training the model to reconstruct the missing parts. In these MAEs, the encoder is replaced with vision transformers (ViTs) [13], and the decoder is a lightweight transformer designed for reconstruction. MAEs have demonstrated performance that, in some instances, outperforms state-of-the-art contrastive learning methods with ViTs, highlighting the efficacy of self-supervised learning in visual domains [14]. Despite these advancements, exploring foundation models within materials science remains limited. Applying self-supervised learning to materials offers the potential to overcome data scarcity and enhance model generalizability across different tasks and material systems. Therefore, in this work, we propose a foundation model specifically designed for composite materials, termed the material masked autoencoder (MMAE), to validate the feasibility of this approach.
We generated a dataset comprising 100,000 grayscale images of short-fiber composites to facilitate self-supervised learning. By pre-training the MMAE on this dataset, the model learns robust and representative embeddings of the composite microstructures without reliance on labeled data. We anticipate that these embeddings capture essential features of the microstructures, making them highly informative for downstream tasks. To evaluate the effectiveness of the pre-trained MMAE, we employed direct numerical simulation (DNS) to generate datasets of homogenized stiffness for two types of composites: short-fiber composites and circular inclusion composites. We then fine-tuned the pre-trained models on significantly smaller labeled datasets to predict the homogenized stiffness of composites. Our results demonstrate that the MMAE achieves high accuracy, with a coefficient of determination (R2) reaching as high as 0.959 and consistently exceeding 0.91 even when trained on limited data. This performance highlights the potential of the MMAE as a foundation model for composite materials and paves the way for more efficient and cost-effective materials design and analysis.
2 Results and Discussion
2.1 Pre-trained MMAE reconstruction performance
As shown in Fig. 1, the pre-trained MMAE accurately reconstructs composite microstructure instances, including both short-fiber and circular inclusion composites, despite the severe masking employed during training. Notably, although the circular inclusion composites were not included in the pre-training dataset, the MMAE could still reconstruct them effectively. These results demonstrate the MMAE’s robustness in learning representative microstructural features and its capacity to generalize to unseen microstructures, indicating its potential for effective transfer learning in downstream tasks.
Figure 1: Reconstruction of unseen images using the pre-trained MMAE with a masking ratio of 85%. Each triplet includes the ground-truth image (left), the masked image (middle), and the reconstructed image (right). Instances #1 to #4 correspond to circular inclusion composites, while instances #5 to #8 correspond to short-fiber composites.
2.2 Transfer learning performance
To evaluate the effectiveness of the embeddings learned by the pre-trained MMAE, we conducted transfer learning experiments on three downstream regression tasks aimed at predicting the homogenized stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$). We utilized two datasets for these tasks: the short-fiber composite dataset and the circular inclusion composite dataset.
2.2.1 Linear probing
To assess the utility of the latent features extracted by the MMAE, we performed linear probing by training a linear regression model on top of the fixed pre-trained embeddings to predict the homogenized stiffness components. We investigated the impact of the masking ratio used during pre-training on the linear probing performance for both the short-fiber and circular inclusion composite datasets.
Fig. 2 presents the linear probing performance on the short-fiber composite dataset. The results indicate that the validation set R2 scores decrease significantly as the masking ratio decreases from 35% to 15%, suggesting that very low masking ratios during pre-training may reduce the informativeness of the latent features for linear probing. However, the performance remains relatively stable and satisfactory when the masking ratio exceeds 45%. Notably, the highest validation set R2 score of 0.93 is achieved at a masking ratio of 85%, indicating that higher masking ratios can lead to more effective feature learning during pre-training. These findings highlight the importance of selecting an appropriate masking ratio to ensure the extraction of meaningful representations that generalize well to downstream tasks.
Fig. 3 presents the linear probing results on the circular inclusion composite dataset. Despite the MMAE being pre-trained on short-fiber composites, the latent features are still effective for predicting the properties of circular inclusion composites. Similar to the short-fiber composite results, an extremely low masking ratio during pre-training adversely affects performance. However, when the masking ratio exceeds 25%, the performance generally remains high, with average validation set R2 scores exceeding 0.975. The slight dip in performance at a masking ratio of 75% suggests some sensitivity to the masking ratio, but overall, the results demonstrate the generality of the learned embeddings across different composite types.
Figure 3: Linear probing performance on the circular inclusion composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
2.2.2 End-to-end fine-tuning
To evaluate the model’s potential to capture nonlinear relationships in the data, we performed end-to-end fine-tuning of the MMAE on the downstream tasks. Unlike linear probing, end-to-end fine-tuning updates all the parameters of the MMAE during training on the downstream tasks, allowing the model to adapt fully to the specific task.
As shown in Fig. 4, the end-to-end fine-tuning performance on the short-fiber composite dataset is relatively insensitive to the masking ratio used during pre-training. In contrast to linear probing, where lower masking ratios led to decreased performance, the end-to-end fine-tuning consistently achieves high average validation set R2 scores of approximately 0.958 across all masking ratios. This result suggests that fine-tuning can effectively adapt the model to the downstream task, regardless of the initial quality of the pre-trained embeddings. The model’s capacity to learn task-specific features during fine-tuning compensates for any deficiencies in the initial representations.
Similarly, Fig. 5 shows the end-to-end fine-tuning performance on the circular inclusion composite dataset. Consistent with the short-fiber composite dataset results, the performance is relatively unaffected by the masking ratio during pre-training. The model achieves validation set R2 scores exceeding 0.99 across all masking ratios, indicating excellent predictive capabilities after fine-tuning. These results underscore the MMAE’s flexibility and capacity to adapt to different composites through fine-tuning.
Figure 5: End-to-end fine-tuning performance on the circular inclusion composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
2.2.3 Partial fine-tuning
Fig. 6 illustrates the performance on the short-fiber composite dataset as a function of the number of blocks fine-tuned. Here, tuning 0 blocks corresponds to linear probing, and tuning all 12 blocks corresponds to end-to-end fine-tuning. The results show that the accuracy improves as more blocks are fine-tuned, indicating better predictive performance. Notably, the performance increases significantly when the last two blocks are fine-tuned and then begins to plateau, suggesting that fine-tuning even a tiny portion of the model’s parameters can capture the necessary task-specific features.
2.2.4 Model explainability
As shown in Fig 7, during linear probing, the saliency maps indicate that the model assigns higher importance predominantly to the matrix regions of the microstructures when predicting all stiffness components. This suggests that the model, constrained by linear regression and with frozen encoder weights, relies mainly on the matrix phase to infer the homogenized stiffness. The emphasis on the matrix implies that the model is not fully capturing the contributions of the fibers and their interactions with the matrix, which are critical to the composite’s mechanical properties. This limitation potentially hinders the model’s predictive accuracy.
In contrast, Fig 8 shows that after end-to-end fine-tuning, the saliency maps highlight the matrix-fiber interfaces more prominently. This shift in focus indicates that the fine-tuned model has learned to capture the nonlinear relationships and critical interactions at the interfaces that significantly influence the composite’s overall stiffness. By paying attention to these regions, the model effectively accounts for factors such as fiber orientation and aspect ratio, essential for accurately predicting the homogenized stiffness components.
2.2.5 Impact of dataset size in transfer learning
To evaluate the influence of dataset size during transfer learning, we selected the MMAE pre-trained with a masking ratio of 85% and varied the number of instances in the fine-tuning dataset. As shown in Fig. 9, we observed that the model’s average validation set R2 score is relatively low when the dataset is very small. However, even with a dataset containing only 500 instances, the model achieves a validation set R2 score above 0.86, demonstrating acceptable performance. As the number of data instances increases, the performance improves, but with diminishing marginal returns. The average R2 score converges when data instances reach approximately 2,500, indicating that adding more data beyond this point yields minimal performance gains.
These findings open avenues for accelerating materials design and analysis, particularly in data-scarce environments. Building upon our results, future research could focus on extending the MMAE framework to three-dimensional composite materials, polycrystalline materials, and other heterogeneous material systems. Such extensions would test the model’s ability to handle more complex geometries and microstructural features inherent in real-world materials.
Additionally, exploring downstream tasks involving more complex, nonlinear material behaviors could further demonstrate the versatility of the MMAE. For instance, integrating the microstructural features extracted by the MMAE with deep material networks [17] could enable more effective extrapolation and prediction of material nonlinearities. This approach would facilitate modeling phenomena such as plasticity, damage, and fracture, which are critical for designing materials with tailored mechanical responses. By advancing the capability of foundation models like the MMAE to encompass a broader range of material behaviors and structures, we can accelerate the discovery and optimization of advanced materials, ultimately contributing to significant advancements in computational materials science and engineering.
In conclusion, our study demonstrates that the MMAE effectively captures essential features of composite microstructures through self-supervised pre-training, enabling high accuracy in predicting homogenized stiffness components with minimal labeled data and computational resources. The combination of robust feature learning, adaptability through fine-tuning, and efficiency in data utilization positions the MMAE as a powerful foundation model for composite materials. These findings suggest that similar approaches could be extended to more complex material systems, further advancing the field of machine learning in materials science.
@@ 3.1 Masked autoencoders
The MMAE employed in this study was primarily based on the MAE architecture proposed by He et al. [12]. As depicted in Fig. 10, the MMAE consists of an encoder-decoder structure, where the encoder is a standard ViT [13] with 12 transformer blocks, and the decoder comprised 8 transformer blocks. Each composite image, originally of size $224\times 224$ pixels, was partitioned into 196 non-overlapping patches through a process known as patchification. Each patch was of size $16\times 16$ grayscale pixels. These patches were then flattened and embedded into tokens suitable for input into the transformer encoder.
@@
3.2 Transfer learning
@@
To evaluate the effectiveness of the embeddings learned by the pre-trained MMAE for predicting homogenized material properties, we employed transfer learning techniques utilizing the [cls] token output from the encoder as a global representation of the input microstructure. We explored linear probing and fine-tuning approaches, as illustrated in Fig. 11.
@@ 3.3.1 Generation of microtrsturce
For the short-fiber composites, we considered representative volume elements (RVEs) containing elliptical inclusions (fibers) dispersed within a matrix. The microstructures were defined within a three-dimensional descriptor space characterized by:
@@
We generated two datasets using Latin hypercube sampling [19] for the short-fiber composites. A large dataset comprising 100,000 grayscale images of short-fiber composite microstructures was used for self-supervised pre-training of the MMAE. This dataset captured various fiber arrangements, aspect ratios, and volume fractions, enabling the MMAE to learn robust and generalizable features from diverse microstructures. An additional dataset of 5,000 short-fiber composite microstructures was generated for the downstream regression tasks. DNS was performed on these microstructures to compute their homogenized stiffness components. The dataset was divided into training and validation sets for these tasks with an 80/20 split.
For the circular inclusion composites, we generated 5,000 microstructures by uniformly sampling the volume fraction $v_{f}$ within the range of 10% to 40%. The inclusions were circular, and their sizes were fixed. The inclusions varied according to the specified volume fractions to maintain the desired $v_{f}$. This dataset evaluated the MMAE’s transfer learning performance on a different composite system with distinct microstructural characteristics.
3.3.2 Direct numerical simulation
To obtain the homogenized stiffness components of the generated microstructures, we performed DNS using the RVE module available in LS-DYNA [20]. The simulations were conducted under periodic boundary conditions to capture the effective mechanical response of the composites accurately.
@@ 3.3.2 Direct numerical simulation
  • Matrix: Young’s modulus $E=100$ GPa, Poisson’s ratio $\nu=0.30$.
  • Short-fiber/Circular-inclusion: Young’s modulus $E=500$ GPa, Poisson’s ratio $\nu=0.19$.
We assumed linear elastic material behavior for both phases to compute the homogenized stiffness components. The focus was on predicting the homogenized stiffness components $\bar{C}_{1111}$, $\bar{C}_{2222}$, and $\bar{C}_{1212}$, which are critical for characterizing the elastic properties of the composites.
The DNS computations provided the necessary labels (homogenized stiffness components) for training and validation during transfer learning. Using consistent material properties and simulation conditions, we ensured the datasets were suitable for assessing the MMAE’s ability to learn and generalize effective material representations.

### v2

@@ Abstract
The rapid advancement of machine learning has unlocked numerous opportunities for materials science, particularly in accelerating the design and analysis of materials. However, a significant challenge lies in the scarcity and high cost of obtaining high-quality materials datasets. While foundation models pre-trained on large datasets have excelled in fields like natural language processing by leveraging latent features through transfer learning, their application in materials science remains limited. Here, we present a foundation model specifically designed for composite materials. Pre-trained on a dataset of short-fiber composites to learn robust latent features, the model accurately predicts homogenized stiffness during transfer learning, even with limited training data. Additionally, our model effectively predicts the material’s nonlinear behavior by transferring these learned features to an Interaction-based Material Network, which is a constitutive surrogate model. These results demonstrate the potential of our foundation model to capture complex material behaviors. Our findings validate the feasibility and effectiveness of foundation models in composite materials. We anticipate extending this approach to more complex three-dimensional composite materials, polycrystalline materials, and beyond. Moreover, this framework enables high-accuracy predictions even when experimental data are scarce, paving the way for more efficient and cost-effective materials design and analysis.
Keywords Foundation model  $\cdot$ Composite material  $\cdot$ Transfer learning  $\cdot$ Self-supervised learning  $\cdot$ Masked autoencoder
@@ 1 Introduction
The concept of foundation models has also been successfully applied to computer vision. Masked autoencoders (MAEs), introduced by He et al. [12], employ a strategy analogous to BERT by masking portions of the input image and training the model to reconstruct the missing parts. In these MAEs, the encoder is replaced with vision transformers (ViTs) [13], and the decoder is a lightweight transformer designed for reconstruction. MAEs have demonstrated performance that, in some instances, outperforms state-of-the-art contrastive learning methods with ViTs, highlighting the efficacy of self-supervised learning in visual domains [14].
Despite these advancements, exploring foundation models within materials science remains limited. Applying self-supervised learning to materials offers the potential to overcome data scarcity and enhance model generalizability across different tasks and material systems. Therefore, in this work, we propose a foundation model specifically designed for composite materials, termed the material masked autoencoder (MMAE), to validate the feasibility of this approach.
To validate the MMAE, we pre-trained the model on a large dataset of composite microstructures and evaluated its performance on two downstream tasks. First, we applied transfer learning to predict the homogenized stiffness components of composite materials, achieving high accuracy even with limited labeled data. Second, we extended the transfer learning framework to predict Interaction-based Material Network (IMN) parameters, enabling the extrapolation of the nonlinear behavior of composite materials [15, 16]. These results demonstrate that the MMAE effectively adapts to various downstream tasks in materials modeling and analysis.
The remainder of the paper is organized as follows: Section 2 outlines the methodology, including data generation and model architecture. Section 3 presents the results and discussion of our experiments. Finally, Section 4 provides concluding remarks and future directions.
@@ 2.1 Masked autoencoders
The MMAE employed in this study was primarily based on the MAE architecture proposed by He et al. [12]. As depicted in Figure 1, the MMAE consists of an encoder-decoder structure, where the encoder is a ViT [13] with a patch size of $16\times 16$ pixels, embedding dimension of 256, 12 transformer blocks, and 4 attention heads per block. A linear layer bridges the encoder and decoder, projecting the encoder’s output embeddings to the appropriate dimension for the decoder. The decoder, tailored for reconstruction, comprises 8 transformer blocks with an embedding dimension of 128 and 16 attention heads per block.
Each composite image, originally of size $224\times 224$ grayscale pixels(1 channel), was divided into 196 non-overlapping patches, each measuring $16\times 16$ pixels. These patches were processed using a 2D convolutional layer (Conv2D) to generate token embeddings, which were subsequently sent to the MMAE encoder.
@@
2.2 Transfer learning for homogenized stiffness components
@@
To evaluate the effectiveness of the embeddings learned by the pre-trained MMAE for predicting homogenized material properties, we employed transfer learning techniques utilizing the [cls] token output from the encoder as a global representation of the input microstructure. We explored linear probing and fine-tuning approaches, as illustrated in Figure 2.
@@ 2.2.2 Fine-tuning
2.3 Transfer learning for IMN parameters
This framework employs transfer learning by fine-tuning the pre-trained MMAE to predict the IMN parameters. The MMAE encodes microstructural information into embeddings, which are passed through a linear projection layer to generate the IMN parameters. During online prediction, the IMN utilizes these parameters to extrapolate nonlinear material behavior, enabling efficient and accurate modeling of complex materials. The schematic of the proposed FM-IMN framework is shown in Figure 3.
Figure 3: Framework of the FM-IMN illustrated for a 2D two-phase RVE. $\mathbf{I}_{s}$ represents the image representation, while $\mathbf{C}^{p1}_{s}$, $\mathbf{C}^{p2}_{s}$, and $\overline{\mathbf{C}}^{\text{DNS}}_{s}$ denote stiffness matrices generated by DNS.
2.3.1 End-to-end offline fine-tuning
The IMN is an efficient physics-based constitutive surrogate model used for multiscale modeling of materials [15, 16, 17]. A notable advantage of the IMN is its ability to be trained using only linear elastic data while achieving accurate nonlinear extrapolation during online prediction. This capability stems from the IMN architecture, which incorporates de-homogenization and homogenization processes during online prediction, enabling it to model complex material behaviors without requiring extensive nonlinear datasets.
The IMN comprises two main components: material nodes and the material network. Specifically, an IMN with a layer depth of $N$ consists of $2^{N}$ material nodes and a perfect binary tree structure called the material network, which contains $12^{N}-$ 1 tree nodes. Each material node represents a material constituent phase and is associated with a trainable parameter $w$ encoding the volume fraction of that phase. Each tree node in the material network has a trainable parameter $\theta$, which represents the stress-equilibrium direction. The collection of all trainable parameters forms $\mathcal{F}$, with a total of $12^{N+1}-$ 1 parameters.
To leverage the microstructural features extracted by the MMAE, we used the [CLS] token output from the encoder, representing a latent embedding of the input microstructure image $I_{s}$. This [CLS] token was passed through a linear projection layer to produce the flattened IMN parameters $\mathcal{F}_{s}$ for each sample $s$. Subsequently, we directly used these parameters $\mathcal{F}_{s}$ to parameterize the IMN, which predicted the homogenized stiffness matrix $\overline{\mathbf{C}}^{\text{MMAE-IMN}}_{s}$, given the constituent stiffness matrices $\mathbf{C}^{p1}_{s}$ and $\mathbf{C}^{p2}_{s}$.
End-to-end fine-tuning was performed to jointly optimize the MMAE encoder and the linear projection layer parameters. To quantify the error between the predicted homogenized stiffness $\overline{\mathbf{C}}^{\text{MMAE-IMN}}_{s}$ and the ground truth $\overline{\mathbf{C}}^{\text{DNS}}_{s}$, we computed the mean squared error (MSE) using the Frobenius norm:
$\text{ERR}=\frac{1}{N_{\text{dataset}}}\sum_{s=1}^{N_{\text{dataset}}}\frac{\left\|\overline{\mathbf{C}}^{\text{DNS}}_{s}-\overline{\mathbf{C}}^{\text{MMAE-IMN}}_{s}(I_{s},\mathbf{C}^{p1}_{s},\mathbf{C}^{p2}_{s})\right\|_{\text{F}}}{\left\|\overline{\mathbf{C}}^{\text{DNS}}_{s}\right\|_{\text{F}}},$ (1)
where $N_{\text{dataset}}$ is the total number of samples, and $\|\cdot\|_{\text{F}}$ denotes the Frobenius norm. We fine-tuned the MMAE encoder and the linear projection layer to accurately predict the homogenized stiffness matrices based on the microstructure images and the constituent material stiffness matrices by minimizing this error.
2.3.2 Online prediction
After training, the MMAE encoder and linear projection layer can infer the IMN parameters $\mathcal{F}$ for any given RVE microstructure image $I$. The inferred IMN, leveraging its online prediction algorithm and the constitutive laws of the constituent phases, performs nonlinear extrapolation to predict the material’s response under diverse loading conditions [15, 16].
By integrating the MMAE’s capability to extract detailed microstructural embeddings with the IMN robust nonlinear extrapolation, this framework provides an efficient and versatile tool for nonlinear material modeling.
@@ 2.4.1 Generation of microtrsturce
For the short-fiber composites, we considered representative volume element (RVE) containing elliptical inclusions (fibers) dispersed within a matrix. The microstructures were defined within a three-dimensional descriptor space characterized by:
@@
We generated three datasets of RVEs using Latin hypercube sampling [19] for short-fiber composites:
  1. 1.
    Pre-training dataset: A large dataset comprising 100,000 grayscale images of short-fiber composite microstructures was generated for self-supervised pre-training of the MMAE. This dataset captures a wide range of fiber arrangements, aspect ratios, and volume fractions, enabling the MMAE to learn robust and generalizable features from diverse microstructures.
  2. 2.
    Homogenized stiffness dataset: A dataset of 5,000 microstructures was generated for predicting homogenized stiffness components.
  3. 3.
    IMN dataset: A smaller dataset of 300 grayscale images was created for IMN transfer learning to predict nonlinear material behavior.
For the circular-inclusion composites, we generated 5,000 microstructures by uniformly sampling the volume fraction $v_{f}$ within the range of 10% to 40%. The inclusions were circular, and their sizes were fixed. The number of inclusions varied according to the specified volume fractions to maintain the desired $v_{f}$. This dataset was utilized to assess the MMAE’s transfer learning performance on a distinct composite system characterized by different microstructural features. This dataset was used for the downstream task of predicting homogenized stiffness components.
2.4.2 Dataset for homogenized stiffness conponents transfer learning
To obtain the homogenized stiffness components of the generated microstructures, we performed DNS using the RVE module available in LS-DYNA [20]. The simulations were conducted under periodic boundary conditions to accurately capture the effective mechanical response of the composites.
@@ 2.4.2 Dataset for homogenized stiffness conponents transfer learning
  • Matrix: Young’s modulus $E=100$ , Poisson’s ratio $\nu=0.30$.
  • Short-fiber/Circular-inclusion: Young’s modulus $E=500$ , Poisson’s ratio $\nu=0.19$.
Assuming linear elastic material behavior, we focused on predicting homogenized stiffness components: $\bar{C}_{1111}$, $\bar{C}_{2222}$, and $\bar{C}_{1212}$. The DNS computations provided the labels required for training and validation during transfer learning. The dataset was split 80/20 into training and validation subsets.
2.4.3 Dataset for IMN transfer learning
For IMN transfer learning, we used a dataset comprising 300 RVEs of short-fiber composite microstructures. To account for material property variability, we augmented the dataset by sampling 300 unique pairs of stiffness matrices $\mathbf{C}^{p1}_{s}$ and $\mathbf{C}^{p2}_{s}$ for each RVE, representing the matrix (Phase 1) and fiber (Phase 2) materials, respectively. This sampling methodology follows the approach outlined in [5].
For each pair of material properties, DNS was performed using LS-DYNA to compute the corresponding homogenized stiffness matrix $\overline{\mathbf{C}}^{\text{DNS}}_{s}$, resulting in a total of $N_{\text{dataset}}=90{,}000$ instances. The dataset was split into training and validation subsets with an 80/20 ratio, allocating 240 RVEs for training and 60 RVEs for validation.
@@ 2.5 Saliency map
3 Results and Discussion
3.1 Pre-trained MMAE reconstruction performance
As shown in Figure 4, the pre-trained MMAE accurately reconstructs composite microstructure instances, including both short-fiber and circular-inclusion composites, despite the severe masking employed during training. Notably, although the circular-inclusion composites were not included in the pre-training dataset, the MMAE could still reconstruct them effectively. These results demonstrate the MMAE’s robustness in learning representative microstructural features and its capacity to generalize to unseen microstructures, indicating its potential for effective transfer learning in downstream tasks.
Figure 4: Reconstruction of unseen images using the pre-trained MMAE with a masking ratio of 85%. Each triplet includes the ground-truth image (left), the masked image (middle), and the reconstructed image (right). Instances #1 to #4 correspond to circular-inclusion composites, while instances #5 to #8 correspond to short-fiber composites.
3.2 Transfer learning for homogenized stiffness components
To evaluate the effectiveness of the embeddings learned by the pre-trained MMAE, we conducted transfer learning experiments on three downstream regression tasks aimed at predicting the homogenized stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$). We utilized two datasets for these tasks: the short-fiber composite dataset and the circular-inclusion composite dataset.
3.2.1 Linear probing
To assess the utility of the latent features extracted by the MMAE, we performed linear probing by training a linear regression model on top of the fixed pre-trained embeddings to predict the homogenized stiffness components. We investigated the impact of the masking ratio used during pre-training on the linear probing performance for both the short-fiber and circular-inclusion composite datasets.
Figure 5 presents the linear probing performance on the short-fiber composite dataset. The results indicate that the validation set R2 scores decrease significantly as the masking ratio decreases from 35% to 15%, suggesting that very low masking ratios during pre-training may reduce the informativeness of the latent features for linear probing. Nevertheless, the performance remains stable and satisfactory even when the masking ratio exceeds 45%. Notably, the highest validation set R2 score of 0.93 is achieved at a masking ratio of 85%, indicating that higher masking ratios can lead to more effective feature learning during pre-training. These findings highlight the importance of selecting an appropriate masking ratio to ensure the extraction of meaningful representations that generalize well to downstream tasks.
Figure 6 presents the linear probing results on the circular-inclusion composite dataset. Despite the MMAE being pre-trained on short-fiber composites, the latent features are still effective for predicting the properties of circular-inclusion composites. Similar to the short-fiber composite results, an extremely low masking ratio during pre-training adversely affects performance. However, when the masking ratio exceeds 25%, the performance generally remains high, with average validation set R2 scores exceeding 0.975. The slight dip in performance at a masking ratio of 75% suggests some sensitivity to the masking ratio, but overall, the results demonstrate the generality of the learned embeddings across different composite types.
Figure 6: Linear probing performance on the circular-inclusion composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
3.2.2 End-to-end fine-tuning
We employed end-to-end fine-tuning on downstream tasks to evaluate the MMAE’s capacity for modeling nonlinear relationships in the data. Unlike linear probing, which updates only a subset of parameters, end-to-end fine-tuning optimizes the entire parameter set, allowing the model to adapt comprehensively to task-specific challenges.
As shown in Figure 7, the end-to-end fine-tuning performance on the short-fiber composite dataset is relatively insensitive to the masking ratio used during pre-training. In contrast to linear probing, where lower masking ratios led to decreased performance, the end-to-end fine-tuning consistently achieves high average validation set R2 scores of approximately 0.958 across all masking ratios. This result suggests that fine-tuning can effectively adapt the model to the downstream task, regardless of the initial quality of the pre-trained embeddings.
Similarly, Figure 8 shows the end-to-end fine-tuning performance on the circular-inclusion composite dataset. Consistent with the short-fiber composite dataset results, the performance is relatively unaffected by the masking ratio during pre-training. The model achieves validation set R2 scores exceeding 0.99 across all masking ratios, indicating excellent predictive capabilities after fine-tuning.
Figure 8: End-to-end fine-tuning performance on the circular-inclusion composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
3.2.3 Partial fine-tuning
Figure 9 illustrates the performance on the short-fiber composite dataset as a function of the number of blocks fine-tuned. Here, tuning 0 blocks corresponds to linear probing, and tuning all 12 blocks corresponds to end-to-end fine-tuning. The results show that the accuracy improves as more blocks are fine-tuned, indicating better predictive performance. Notably, the performance increases significantly when the last two blocks are fine-tuned and then begins to plateau, suggesting that fine-tuning even a tiny portion of the model’s parameters can capture the necessary task-specific features.
3.2.4 Model explainability
As shown in Figure 10, during linear probing, the saliency maps indicate that the model assigns higher importance predominantly to the matrix regions of the microstructures when predicting all stiffness components. This suggests that the model, constrained by linear regression and with frozen encoder weights, relies mainly on the matrix phase to infer the homogenized stiffness components. The emphasis on the matrix implies that the model is not fully capturing the contributions of the fibers and their interactions with the matrix, which are critical to the composite’s mechanical properties. This limitation potentially hinders the model’s predictive accuracy.
In contrast, Figure 11 shows that after end-to-end fine-tuning, the saliency maps highlight the matrix-fiber interfaces more prominently. This shift in focus indicates that the fine-tuned model has learned to capture the nonlinear relationships and critical interactions at the interfaces that significantly influence the composite’s overall stiffness. By paying attention to these regions, the model effectively accounts for factors such as fiber orientation and aspect ratio, essential for accurately predicting the homogenized stiffness components.
3.2.5 Impact of dataset size in transfer learning
To evaluate the influence of dataset size during transfer learning, we selected the MMAE pre-trained with a masking ratio of 85% and varied the number of instances in the fine-tuning dataset. As shown in Figure 12, we observed that the model’s average validation set R2 score is relatively low when the dataset is very small. However, even with a dataset containing only 1000 instances, the model achieves a validation set R2 score above 0.91, demonstrating acceptable performance. As the number of data instances increases, the performance improves, but with diminishing marginal returns. The average R2 score converges when data instances reach approximately 2,500, indicating that adding more data beyond this point yields minimal performance gains.
3.3 Transfer learning for IMN parameters
To explore the effectiveness of the pre-trained MMAE for more complex downstream tasks, we optimized the IMN parameters to predict nonlinear composite behavior. This process involved two key stages: offline training through end-to-end fine-tuning and online prediction. The specific details of these stages are outlined below.
3.3.1 Offline training
In the offline training phase, we employed MMAE models pre-trained with varying masking ratios to investigate their impact on transfer learning performance in the short-fiber composite dataset. Specifically, we performed end-to-end fine-tuning to enhance the latent features for predicting IMN parameters, focusing on how the pre-training masking ratios influenced the fine-tuning process.
As shown in Figure 13, the validation error remained relatively stable, ranging from 6% to 7% across all tested masking ratios for the short-fiber composite dataset. Notably, the lowest validation error occurred at a masking ratio of 50%, which was selected for subsequent online prediction analyses.
Figure 13: Validation error during offline training using MMAE models pre-trained with different masking ratios.
3.3.2 Online prediction
In the online prediction phase, the MMAE–IMN framework was evaluated on four previously unseen RVEs subjected to complex loading conditions. Despite being trained on an offline dataset comprising only 300 RVEs, the pre-trained MMAE demonstrated robust generalization to these new microstructures, showcasing its capability for effective feature extraction. As shown in Figure 14, the predicted stress–strain curves closely align with the DNS results, with the corresponding microstructures depicted as insets. These findings highlight the MMAE–IMN framework’s ability to accurately extract microstructural features and extrapolate nonlinear mechanical behavior, even with a limited fine-tuning dataset.
Figure 14: Predicted stress-strain curves for four unseen RVEs under complex loading conditions. Insets show the corresponding microstructures, illustrating the phase distributions (blue: Phase 1, red: Phase 2).
Table 1 summarizes the material properties of Phase 1 and Phase 2. Phase 1 is modeled with elastoplastic isotropic hardening, whereas Phase 2 is treated purely as an elastic phase.
Table 1: Material properties for Phase 1 and Phase 2.
Young modulus $E$ Poisson ratio $\nu$ Yield stress $\sigma_{y}$ Tangent modulus
Phase 1 100 GPa 0.3 0.1 GPa 5 GPa
Phase 2 500 GPa 0.19
3.4 Summary
The downstream tasks demonstrate that MMAE’s self-supervised pre-training is effective across diverse data-scarce scenarios. For linear probing, the pre-trained embeddings facilitate the extraction of generalizable latent representations of composite microstructures, maintaining robustness even under data-scarce conditions. For more demanding tasks, such as IMN parameter calibration, end-to-end fine-tuning enhances the model’s adaptability, enabling it to meet specific task requirements with high precision.
Saliency analyses reveal that fine-tuning refines the model’s focus on critical fiber–matrix interfaces, improving predictive accuracy for homogenized stiffness components. These findings underscore the utility of self-supervised pre-training as a robust foundation for downstream tasks, while highlighting the importance of tailored adaptations to achieve accurate and efficient material property predictions, particularly in data-limited contexts.
4 Conclusion
This work proposed the MMAE, a foundation model tailored for composite materials, addressing the persistent challenges of data scarcity and transferability in materials science. Through self-supervised pre-training on a large dataset of short-fiber composite microstructures, the MMAE effectively learned robust and representative latent features, providing a solid foundation for diverse downstream applications.
In the first transfer learning scenario, we fine-tuned the pre-trained MMAE to predict the homogenized stiffness of composite materials, achieving an impressive $R^{2}$ score of up to 0.959, even with limited labeled data. This result highlights the model’s effectiveness in overcoming data-scarcity challenges. In the second scenario, the MMAE was fine-tuned to infer IMN parameters, enabling precise prediction of composite material responses and demonstrating its adaptability to tasks requiring the modeling of intricate material behaviors.
These findings highlight the transformative potential of foundation models in materials science, offering scalable and efficient solutions for property prediction and microstructural analysis. Future work will explore extending the MMAE framework to more complex systems, such as three-dimensional composites and polycrystalline materials, as well as incorporating experimental datasets to enhance its applicability in real-world scenarios.